OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = 7, an = an-1 + 6 13, 19, 25, 31, 37, 43 0, 6, 12, 18, 24, 30 7, 6, 12, 18, 24, 30 7, 13, 19, 25, 31, 37

OpenStudy (johnweldon1993):

Well we know that a1 = 7 so \[\large a2 = a_{2 - 1} + 6\] meaning \[\large a2 = a_{1} + 6\] We are given that a1 = 7 so \[\large a2 = 7 + 6 = 13\]